You have helped me a few times and I greatly appreciate all I can get, here is some more questions I am stuck on
1. f(x)=2x^2-x+7;g(x)=x-1 find (f-g)(x)
2. f(x)=x^2+3;g(x)=+4; find (fog)(-2)
3. if f(x)=2x-4, find f^-1(x)
4. using the compound interest formula A=P(1+r/n)^nt, find A when P=$1000, r=6%, t=5 years and n=4

Answer by stanbon(48535)

1. f(x)=2x^2-x+7;g(x)=x-1 find
(f-g)(x) = f(x) - g(x = (2x^2 -x +7) - (x - 1) = 2x^2 -2x + 8
--------------------------------------

2. f(x)=x^2+3;g(x)=+4; find
(fog)(-2) = f[g(-2)] = f[4] = 16+3 = 19
------------------------------------
3. if f(x)=2x-4, find f^-1(x)
x = 2y - 4
2y = x+4
y = (1/2)x + 2
-------------------------
4. using the compound interest formula A=P(1+r/n)^nt, find A when P=$1000, r=6%, t=5 years and n=4
A(5) = 1000(1 + 0.06/4)^(4*5)
A(5) = 1000(1.015)^20
A(5) = 1000(1.346855007)
A(5) = $1346.86
=====================
Cheers,
Stan H.